OpenStudy (anonymous):

The average number of 7 numbers is 38. When an eighth number is added, the average increases by1. Find the eighth number

OpenStudy (anonymous):

so the sum of these 7 number is 7*38=266 the eighth number will increase average by 1 so the new sum is 8*39=312 so the eighth no is 312-266=46

OpenStudy (anonymous):

Many thanks